A communications tower is supported by four guy wires. The tower is 155 m tall and each guy wire is staked into the ground at a distance of 30 m from the base of the tower. What is the total length of wire used to support the tower?
Answer by Boreal(15235)

Each guy wire is the hypotenuse of a right triangle with legs 155 m and 30 m
sqrt (hypot)= sqrt (155^2+30^2)=sqrt (24925)=157.88 m
There are four of them. If I round only at the end, the amount of wire is 631.5 or 632 m.